ADNAN MUHAMMED

Grade 12,

the sum of the series 1\2+3\4+7\8+15\16+........, is

the sum of the series 1\2+3\4+7\8+15\16+........, is

Grade:11

1 Answers

Arun
6 years ago
For the sum of the series
tn=(2n-1)/2n
  =1-1/2n
Sn=∑1-1/2n =n-∑1/2n (gp series)
    =n-1+2-n
 
Hope it helps
Happy Learning!